# Squeezing information from radio surveys to probe the primordial Universe

### v1

@@ II PNG from 21cm intensity mapping
Figure 1: Schematic illustration of the ranges of radial and transverse wavenumbers that the two survey modes are sensitive to (adapted from [43]).
The presence of local PNG induces a large-scale scale-dependent behaviour on the power spectrum of a tracer [44, 45, 46, 47, 48, 49]. Hence, a LSS survey should be able to probe the largest scales of clustering in order to provide competitive constraints on $f_{\mathrm{NL}}^{\mathrm{loc}}$. Moreover, attaining a deeper comprehension of galaxy formation is essential for establishing constraints on the amplitude of local PNG [50, 51, 52]. For the bispectrum, access to a wide scale range is essential, in order to probe enough squeezed triangles and boost the signal-to-noise ratio. Thus, a single-dish HI IM survey favours a power spectrum analysis, where has been suggested as well in the past as a powerful probe for local PNG [53, 54, 55, 56, 57], outperforming an interferometer survey. The reverse holds for the bispectrum $f_{\mathrm{NL}}^{\mathrm{loc}}$ constraints, where surveys designed for IF mode outshine those in SD mode [58].
III The 21cm bimodal bispectrum
In this Letter, we propose a novel method to push the capabilities of HI IM to tightly constrain the amplitude of local PNG, while excluding the contaminated $k$-modes from the analysis. The key observation lies in the fact that the two functionalities of the radio telescope probe complimentary scales (see Fig. 1). Combining the data from these two can provide access to a very wide scale range with high resolution: from the large and linear scales, as probed by SD mode, up to the very small and nonlinear ones, measured by IF mode. This is optimal for a bispectrum analysis and especially for an observable that has a strong signal on squeezed triangles, like local PNG.
Squeezed configurations correlate different scales, i.e. one side of the triangle is much smaller than the other two ($k_{1}\sim k_{2}\gg k_{3}$). Thus, cross correlating a density field $\delta^{\rm SD}(\bm{k})$, as probed by a single-dish mode survey, with two density fields $\delta^{\rm IF}(\bm{k})$ probed by a survey in interferometer mode, we can form an HI IM bispectrum that has access to a very large number of squeezed triangles. We define the bimodal bispectrum, synergising two HI IM surveys in two different operating modes, for an overlapping redshift range and sky area, as
$\langle\delta^{\rm IF}(\bm{k}_{1})\,\delta^{\rm IF}(\bm{k}_{2})\,\delta^{\rm SD}(\bm{k}_{3})\rangle=\\ \,(2\,\pi)^{3}\,\delta_{\rm D}(\bm{k}_{123})\,B^{\rm SD\times IF}(\bm{k}_{1},\bm{k}_{2},\bm{k}_{3})\;,$ (1)
where $\bm{k}_{123}\equiv\bm{k}_{1}+\bm{k}_{2}+\bm{k}_{3}$ and the Dirac $\delta_{\rm D}$ ensures momentum conservation.
The fields $\delta^{\rm SD,IF}$ are fluctuations in the 21cm brightness temperature and their measurements in Fourier space depend on survey window functions $W^{\rm SD,IF}(\bm{k})$. Crucially, SD observations observe the actual sky, so that the fluctuation is a convolution in Fourier space, ${\delta^{\rm SD}(\bm{k})=W^{\rm SD}(\bm{k})\star\delta(\bm{k})}$, whereas IF mode measures a Fourier transform of the sky, leading to a simple product, ${\delta^{\rm IF}(\bm{k})=W^{\rm IF}(\bm{k})\,\delta(\bm{k})}$. Here and below we omit the $z$-dependence for brevity.
@@ III The 21cm bimodal bispectrum
The estimator for measuring the synergy bispectrum between the SD and IF modes of the neutral hydrogen instrument is
$\hat{B}^{\rm SD\times IF}(\bm{k}_{1},\bm{k}_{2},\bm{k}_{3})$ $=\frac{1}{V_{s}\,V_{123}}\sum_{\bm{q}_{1}\in k_{1}}\sum_{\bm{q}_{2}\in k_{2}}\sum_{\bm{q}_{3}\in k_{3}}\delta_{\rm K}(\bm{q}_{123})$
×$\times\,\delta^{\rm IF}(\bm{q}_{1})\,\delta^{\rm IF}(\bm{q}_{2})\,\delta^{\rm SD}(\bm{q}_{3})\;,$ (2)
where $V_{s}$ is the volume of the survey within each redshift bin. The Kronecker delta ensures the formation of ‘fundamental triangles’ with sides $\bm{q}_{i}$, satisfying $\bm{q}_{123}=0$, that fall in the ‘triangular bin’ of width $\Delta k$ defined by the bin centres ($k_{1}$, $k_{2}$, $k_{3}$). The volume in Fourier space of the fundamental triangle bins is, in the thin shell limit $V_{123}=8\,\pi^{2}\,k_{1}\,k_{2}\,k_{3}\,\Delta k^{3}$ [67], where the bin size $\Delta k$ is taken to be the fundamental frequency of the survey, $\Delta k=k_{\rm f}=2\pi/V_{\rm s}^{1/3}$. For the extreme cases of flattened ($k_{1}=k_{2}+k_{3}$) and open triangles ($k_{1}\neq k_{2}+k_{3}$) this expression breaks. Here we consider the analytic corrections for these configurations presented in [68].
The covariance of the bimodal bispectrum estimator is given by
${\sf C}_{ij}\equiv\big{\langle}\delta\hat{B}_{i}\,\delta\hat{B}_{j}\big{\rangle}={\sf C}_{ij}^{\rm G}+{\sf C}_{ij}^{\rm NG}\,,$ (3)
where $\delta\hat{B}_{i}=\hat{B}_{i}-\langle\hat{B}_{i}\rangle$, $i$ and $j$ indices denote the different triangle configuration, i.e. $B_{i}\equiv B(k_{1}^{i},k_{2}^{i},k_{3}^{i})$, and G, NG are the Gaussian and non-Gaussian contributions. In the thin shell limit ($k\gg\Delta k$), the Gaussian part can be written as
${\sf C}_{ij}^{\rm G}=\,\frac{(2\pi)^{6}}{V_{s}\,V_{123}}\,s_{123}\,\delta_{ij}\,P_{\rm HI}^{\rm IF}(\bm{k}_{1})\,P_{\rm HI}^{\rm IF}(\bm{k}_{2})\,P_{\rm HI}^{\rm SD}(\bm{k}_{3})\,,$ (4)
where $s_{123}=6,2,1$ for equilateral, isosceles and scalene triangles respectively. The HI IM power spectrum is the tree-level model, including the instrumental noise and shot noise [58] (see also the Appendix).
The non-Gaussian part of the covariance is a complicated quantity to calculate analytically. However, in the case of squeezed configurations, like those considered in this work, it can be accurately approximated by [69, 68, 70]
${\sf C}_{ij}^{\rm NG}$ $=\frac{2(2\,\pi)^{3}}{V_{s}\,V_{123}^{i}\,V_{123}^{j}}\Big{[}\delta_{k_{1}^{i}k_{1}^{\,j}}\,U(k_{1}^{i},k_{1}^{\,j})$
×$\times B_{\rm HI}(\bm{k}_{1}^{\,j},\bm{k}_{2}^{i},\bm{k}_{3}^{i})\,B_{\rm HI}(\bm{k}_{1}^{i},\bm{k}_{2}^{\,j},\bm{k}_{3}^{\,j})+8\,{\rm perm}\Big{]},$ (5)
where $U(k_{1}^{i},k_{1}^{j})=16\pi^{3}\,k_{2}^{i}\,k_{3}^{i}\,k_{2}^{\,j}\,k_{3}^{\,j}\Delta k^{5}$. For the extreme triangle configurations, we consider the analytic corrections presented in Ref. [68]. For the HI IM bispectrum, $B_{\rm HI}$, the tree-level model is considered [58] (see also the Appendix). A well motivated assumption for the instrumental noise is that it is Gaussian [43]; in this case the non-Gaussian part of the covariance will not contain instrumental noise contributions.
IV Results
In this section, we present forecasts for the HI IM bimodal bispectrum on constraining the amplitude of local PNG. The Fisher information matrix formalism is used, which provides the minimum error on a parameter achieved by a survey setup, while requiring only an analytic model as an input. In a given redshift bin the Fisher matrix is
$F_{\alpha\beta}=\int{\rm d}^{3}\bm{k}_{i}\,\frac{\partial B^{\rm SD\times IF}(\bm{k}_{i})}{\partial\theta_{\alpha}}\,{\sf C}_{ij}^{-1}\,\frac{\partial B^{\rm SD\times IF}(\bm{k}_{j})}{\partial\theta_{\beta}}\,.$ (6)

### v2

@@ 4 Results
All work on constraining $f_{\mathrm{NL}}^{\mathrm{loc}}$ is affected by a degeneracy between $f_{\mathrm{NL}}^{\mathrm{loc}}$ and $b_{\phi}$ [76, 77], where $b_{\phi}$ is the local PNG bias parameter [50, 78]. This compromises the ability of the LSS power spectrum and bispectrum to tightly constrain $f_{\mathrm{NL}}^{\mathrm{loc}}$. PNG bias depends on the halo formation history [79], i.e., on properties beyond total mass. If it is accurately modeled and constrained, e.g. with hydrodynamical simulations, this degeneracy is broken and the power of the LSS bispectrum to constrain $f_{\mathrm{NL}}^{\mathrm{loc}}$ is completely restored [80]. Our focus here is on the relative improvement from a new bimodal synergy, compared to the sum of the standard bispectrum information from 2 surveys – therefore, our results are not affected by the model for $b_{\phi}$, and we use the simplest model that breaks the degeneracy [46, 48, 47].
After marginalization over the free parameters of the model [57, 75], we retrieve the final forecasts on $f_{\mathrm{NL}}^{\mathrm{loc}}$. The two surveys that we consider are with Band 1 of SKA-MID [30], which is a radio telescope in SD mode and has a redshift range of z=0.353.05z=0.35-3.05, and the final version of HIRAX [31] in IF mode with z=0.752.55z=0.75-2.55. We consider only the signal from the overlapping range (z=0.752.55z=0.75-2.55) and sky coverage ($15,000$ $\rm deg^{2}$). For details of the surveys considered and the specifics of the instrumental noise, see Appendix˜A and [75].
The forecasts on $f_{\mathrm{NL}}^{\mathrm{loc}}$ from the synergy between SKA and HIRAX are presented in Fig.˜2, for the conservative and optimistic $k_{\rm max}$ cases. The results when removing the foreground cut, i.e. when assuming perfect foreground cleaning and signal reconstruction, are shown in order to illustrate the limiting capabilities of the surveys. We also present the forecasts from each individual survey, as well as their summed signal (SKA+HIRAX) without cross-correlation, for the same redshift range and squeezed configurations as the bimodal case. The results indicate that the bimodal bispectrum provides a significant improvement on $f_{\mathrm{NL}}^{\mathrm{loc}}$ constraints by a factor of $\sim 4$, compared to those from each individual survey and a factor of $\sim 2-3$ on their summed signal.
The constraints are competitive with current CMB bounds by a factor of $\sim 2$. The bimodal bispectrum achieves this by only using a small subset, $\mathcal{O}(10^{2}-10^{3})$, of the total number of formed triangles. This means that the forecasts could improve further once all the triangles are considered, even achieving $\sigma(f_{\mathrm{NL}}^{\mathrm{loc}})\lesssim 1$, as in the case of dedicated surveys with much larger sky area (e.g. SPHEREx [81] or PUMA [82]). In addition, by including the single-survey information from the non-overlapping redshifts and sky area, the constraints would improve even further and utilise to the fullest the available signal in the SKA-HIRAX synergy.
Figure 3: Normalized non-Gaussian SNR [Eq.˜4.2] as a function of the long- and short-wavelength modes of the squeezed configuration (i.e. $k_{1}=k_{2}=k_{\rm S}$ and $k_{3}=k_{\rm L}$), after integrating over the triangle orientation with the line-of-sight. Upper panels are at redshift $z=1$, lower at $z=2.5$. The left column is the fiducial case, while the right is the case where no radial foregrounds are considered (i.e. no $k_{\parallel{\rm min}}$ cut).
To investigate the effect of the radial foreground cut, we calculate the PNG signal-to-noise ratio (SNR) of the bimodal bispectrum, as
$({\rm S/N})_{\rm PNG}^{2}(z)=\int{\rm d}^{3}\bm{k}\,{B_{\rm PNG}^{\rm SD\times IF}(\bm{k}_{i},z)}\,\left({\sf C}^{\rm G}\right)^{-1}_{ij}\,B_{\rm PNG}^{\rm SD\times IF}(\bm{k}_{j},z)\;,$ (4.2)
where $B_{\rm PNG}=B(f_{\mathrm{NL}}^{\mathrm{loc}}\neq 0)-B(f_{\mathrm{NL}}^{\mathrm{loc}}=0)$. The results are presented in Fig.˜3 as a function of the squeezed triangle’s sides, i.e. the long-wavelength ($k_{3}=k_{\rm L}$) and short-wavelength ($k_{1}=k_{2}=k_{\rm S}$) modes, after integrating over the triangle orientation angles. Shown at redshifts $z=1$ and $z=2.5$, two scenarios are considered: excluding $k$-modes affected by radial foregrounds ($k_{\parallel}<k_{\parallel{\rm min}}$, left panels) and fully recovering these modes (right panels).
At low redshifts ($z=1$), recovering the $k$-modes lost to foreground contamination is essential for shifting part of the PNG signal into the linear regime. The dominant contribution to the local PNG signal lies in bispectrum modes where $k_{1}\sim k_{2}=k_{\rm S}>k_{\rm max}$, i.e. beyond the range of linear theory. When long-wavelength radial modes are lost to the foregrounds, only a small part of the maximal PNG S/N is accessible. However, if these modes are successfully reconstructed, more squeezed triangles fall within the perturbative regime, making them amenable to analytic modeling and significantly enhancing the contribution of low-redshift data to the PNG constraints.
In contrast, at high redshifts ($z=2.5$), where the Universe is more linear, the majority of the PNG signal remains within the validity of perturbation theory, even in the presence of a radial foreground cut. As shown in the lower panels of Fig.˜3, these redshifts already contribute significantly to the total PNG constraints. The tree-level model remains reliable, and a large number of squeezed configurations – particularly those with maximal PNG signal–can still be utilized. While probing smaller scales, i.e. $k_{\rm S}>k_{\rm max}$, would provide access to more configurations, it is not essential for these high-redshift bins to yield strong constraints on $f_{\mathrm{NL}}^{\mathrm{loc}}$.
This explains why, as illustrated in Fig.˜2, reconstructing the lost long-wavelength radial $k$-modes yields a greater improvement in $f_{\mathrm{NL}}^{\mathrm{loc}}$ forecasts than increasing $k_{\rm max}$. Although accessing very large scales is critical for HI intensity mapping, exploring mildly nonlinear regimes offers a complementary strategy to mitigate information loss due to 21cm foregrounds.
To assess the sensitivity of the results to the assumed value of $k_{\parallel{\rm min}}$, we repeat the analysis for $k_{\parallel{\rm min}}=0.01\;\text{$h$/Mpc}$. Increasing $k_{\parallel{\rm min}}$ impacts all surveys and scenarios, with the effect being most significant for the SD mode survey. Specifically, constraints for the SD mode degrade by a factor of $\sim 2.5$ compared to the $k_{\parallel{\rm min}}=0.005\;\text{$h$/Mpc}$ case, while the impact on the IF mode is minimal ($\sim 15\%$). For the bimodal forecasts, the degradation is by a factor of $\sim 2$, and for the summed signal from the SD and IF modes, by a factor of $\sim 1.5$. These results indicate that the advantage of the bimodal approach over the combined signal from the SD and IF surveys remains approximately the same, with a relative improvement by a factor of $\sim 2$.
To evaluate how the ability of the SD survey to probe the large scales influences $f_{\mathrm{NL}}^{\mathrm{loc}}$ constraints, we vary its total integration time and the number of dishes. The results show that even a tenfold reduction in survey capabilities leads to only a a $\sim 10\%$ loss in precision, suggesting that instrumental noise of the SD survey has a limited effect on the constraints. For a more detailed discussion, see Appendix˜A.
5 Conclusions
We have shown how to combine SD and IF observations of HI IM maps in a novel way to give significantly stronger constraints on $f_{\mathrm{NL}}^{\mathrm{loc}}$ than was previously thought possible. In fact, this method will improve on the Planck results by around a factor of 2, using only a few thousand triangles per redshift bin. The bimodal bispectrum improves constraints relative to the sum of the individual surveys by a factor of $\sim 4$, which is equivalent to a much larger SD or IF array.
The key aspect of combining SD and IF modes, is that we identified a scale regime where instrumental noise is small and free from other systematics. This enables exceptionally clean measurements of the squeezed bispectrum, containing sufficient squeezed triangle configurations to impose tight constraints on local PNG. This synergy between SD and IF modes is critical for achieving high-precision results on local PNG.
The surveys considered here (SKA and HIRAX) are not the largest possible, as they are intended to serve as test cases to demonstrate the potential of this technique rather than to enable direct comparisons with forecasts from other current or future surveys. Utilizing larger surveys, such as future HI IM projects (e.g., SKA Phase 2, PUMA), could further tighten constraints, surpassing any current forecasts. Indeed, our results already outperform constraints from current galaxy surveys with comparable fields of view to those used here, such as Euclid [9] and DESI [83], with constraints improving substantially as survey parameters are optimized and information from all available triangle configurations is incorporated.
A substantial part of the constraining power comes from the IF mode signal, implying that the results will improve even further if the IF instrumental noise can be reduced. By contrast, the method is not so sensitive to the SD mode noise. Instead, the SD signal is sensitive to the loss of very large scales $k_{\parallel}<k_{\parallel{\rm min}}$, so that further improvement requires advances in foreground cleaning and long-wavelength $k_{\parallel}$-mode reconstruction. The sensitivity of the results to the assumed value of $k_{\parallel{\rm min}}$ is also examined. While all cases are impacted, the advantage of the bimodal approach over the combined signal from the SD and IF modes remains largely unchanged, highlighting the effectiveness of the proposed approach in constraining local PNG from HI IM surveys.
Long-wavelength modes provide a set of very squeezed triangles that carry a significant amount of signal on local PNG. On the other hand, increasing the number of squeezed triangles by changing $k_{\rm max}$ provides a much less pronounced gain in the $f_{\mathrm{NL}}^{\mathrm{loc}}$ constraints ($\sim 20\%$). The constraints with an optimistic $k_{\rm max}$ are very close ($\sim 15\%$ difference) to those provided by the conservative case, in which no radial foregrounds are considered. This indicates that the local PNG signal lost to foregrounds can be partially recovered, by venturing further into the intermediate regime. Smaller scales have a significant amount of information on PNG, which requires alternative methods to the analytical modelling, e.g. simulation-based inference [18, 19, 20, 21]. We leave this for future work.
The bimodal bispectrum cross-correlates SD large scales with smaller-scale IF modes. We do not expect foregrounds or instrumental systematics to show correlations between long- and short-wavelength $k$-modes, another advantage of the bimodal synergy. The advantages of the bimodal bispectrum are not limited to local PNG studies. Any observable that has signal peaking on squeezed triangles can benefit from the proposed synergy. Moreover, including all triangle configurations into the bimodal analysis, still maintains a significant gain in the constraining power (see Appendix˜A).
Acknowledgments
DK is supported by the Science and Technology Facilities Council (UK) grant number ST/X000931/1. DK, RM and SS acknowledge support for this work from the University of Missouri South African Education Program. RM is supported by the South African Radio Observatory and National Research Foundation (grant no. 75415). SS acknowledges support for this work from NSF-2219212. SS is supported in part by World Premier International Research Center Initiative, MEXT, Japan. SC acknowledges support from the Italian Ministry of University and Research, PRIN 2022 ‘EXSKALIBUR – Euclid-Cross-SKA: Likelihood Inference Building for Universe Research’, from the Italian Ministry of Foreign Affairs and International Cooperation (grant no. ZA23GR03), and from the European Union – Next Generation EU. JF acknowledges support of Fundação para a Ciência e a Tecnologia through the Investigador FCT Contract No. 2020.02633.CEECIND/CP1631/CT0002, the FCT project PTDC/FIS-AST/0054/2021, and the research grants UIDB/04434/2020 and UIDP/04434/2020.
Appendix A HI IM model and forecast details
The surveys considered here are the SKAO333www.skatelescope.org [30], operating in single-dish mode, and HIRAX444hirax.ukzn.ac.za [31], operating in interferometer mode. The survey specifications are presented in Table˜1.
SKAO (Band 1) HIRAX
redshift $^{a} 0.7752.550.775-2.55$
$N_{\rm dish}$ $197$ $1,024$
$D_{\rm dish}$ [m] $15$ 66
$S_{\rm area}$ $\rm{deg}^{2}$] $^{b} 15,00015,000$
$t_{\rm survey}$ [hrs] $10,000$ $17,500$
Table 1: The survey specifications for the two HI intensity mapping (IM) surveys considered in this work. Notes: (a) The full redshift range covered by SKAO Band 1 is z=0.353.05z=0.35-3.05 (see Ref. [30] for details). However, to quantify only the gain of the bi-modal approach, i.e. , we restrict our analysis to the overlapping redshift range, z=0.7752.55z=0.775-2.55, which matches that of HIRAX. (b) While SKAO offers a total sky coverage of $20,000$ deg2, we limit our the analysis to the overlapping sky area of $15,000$ deg2, again to ensure a consistent comparison.
The tree-level HI IM power spectrum and bispectrum model in redshift space for operational modes SD, IF are given by :
$P^{\rm SD,IF}(\bm{k},z)$ $=T_{b}(z)^{2}\Big{[}D_{\text{FOG}}^{P}(\bm{k},z)Z_{1}(\bm{k},z)^{2}P_{m}^{\rm L}(k,z)+P_{\varepsilon}(z)\Big{]}+P_{\rm N}^{\rm SD,IF}(\bm{k},z),$ (A.1)
$B^{\rm SD\times IF}(\bm{k}_{1},\bm{k}_{2},\bm{k}_{3},z)$ $=T_{b}(z)^{3}\,\Bigg{[}D_{\text{FOG}}^{B}(\bm{k}_{1},\bm{k}_{2},\bm{k}_{3},z)$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \times\bigg{[}Z_{1}(\bm{k}_{1},z)Z_{1}(\bm{k}_{2},z)Z_{1}(\bm{k}_{3},z)B_{\rm p}(k_{1},k_{2},k_{3},z)$
$\leavevmode\nobreak\ \leavevmode\nobreak\ +\Big{[}2Z_{1}(\bm{k}_{1},z)Z_{1}(\bm{k}_{2},z)Z_{2}(\bm{k}_{1},\bm{k}_{2},z)P_{m}^{\rm L}(k_{1},z)P_{m}^{\rm L}(k_{2},z)+2\leavevmode\nobreak\ \text{perm}\Big{]}\bigg{]}$
$\leavevmode\nobreak\ \leavevmode\nobreak\ +2P_{\varepsilon\varepsilon_{\delta}}(z)\Big{[}Z_{1}(\bm{k}_{1},z)P_{m}^{\rm L}(k_{1},z)+2\leavevmode\nobreak\ \text{perm}\Big{]}+B_{\varepsilon}(z)\Bigg{]}.$ (A.2)
@@ Appendix A HI IM model and forecast details
The redshift kernels up to second order, which include local PNG, are [90, 91, 9]:
$Z_{1}(\bm{k}_{i},z)$ $=b_{1}(z)+f(z)\mu_{i}^{2}+\frac{f_{\mathrm{NL}}^{\mathrm{loc}}b_{\phi}(z)}{M(k_{i},z)},$ (A.3)
$Z_{2}(\bm{k}_{i},\bm{k}_{j},z)$ $=b_{1}(z)F_{2}(\bm{k}_{i},\bm{k}_{j})+f(z)\mu_{ij}^{2}G_{2}(\bm{k}_{i},\bm{k}_{j})+\frac{b_{2}(z)}{2}+\frac{b_{s^{2}}(z)}{2}S_{2}(\bm{k}_{i},\bm{k}_{j})$
$\leavevmode\nobreak\ \leavevmode\nobreak\ +\frac{f(z)\mu_{ij}k_{ij}}{2}\left[\frac{\mu_{i}}{k_{i}}Z_{1}(\bm{k}_{j},z)+\frac{\mu_{j}}{k_{j}}Z_{1}(\bm{k}_{i},z)\right]$
$\leavevmode\nobreak\ \leavevmode\nobreak\ +f_{\mathrm{NL}}^{\mathrm{loc}}\Bigg{[}\frac{b_{\phi\delta}(z)-b_{\phi}(z)N_{2}(\bm{k}_{j},\bm{k}_{i})}{2M(k_{i},z)}+\frac{b_{\phi\delta}(z)-b_{\phi}(z)N_{2}(\bm{k}_{i},\bm{k}_{j})}{2M(k_{j},z)}\Bigg{]}.$ (A.4)
where $f(z)$ is the linear growth rate, $\mu_{i}=\hat{\bm{k}}_{i}\cdot\hat{\bm{z}}$, with $\hat{\bm{z}}$ being the line-of-sight vector, $\mu_{ij}=(\mu_{i}k_{i}+\mu_{j}k_{j})/k_{ij}$ and $k_{ij}^{2}=(\bm{k}_{i}+\bm{k}_{j})^{2}$. The function $M(k,z)=2c^{2}D(z)\,T(k)\,k^{2}/(3\Omega_{m}H_{0}^{2}\,{g_{\rm dec}})$, contains the linear growth factor $D(z)$ and matter transfer function $T(k)$, and $g_{\rm dec}$ is the Bardeen potential growth factor at decoupling (ensuring that $f_{\mathrm{NL}}^{\mathrm{loc}}$ is in the CMB convention [92]). The kernels $F_{2}(\bm{k}_{i},\bm{k}_{j})$ and $G_{2}(\bm{k}_{i},\bm{k}_{j})$ are the second-order symmetric SPT kernels [93], while $S_{2}(\bm{k}_{1},\bm{k}_{2})=(\hat{\bm{k}}_{1}\cdot\hat{\bm{k}}_{2})^{2}-1/3$ is the tidal kernel [94, 95]. The kernel $N_{2}(\bm{k}_{1},\bm{k}_{2})=(\bm{k}_{1}\cdot\bm{k}_{2})k_{1}^{2}$ encodes the coupling of the PNG potential to the Eulerian-to-Lagrangian displacement field [96, 90]. The bias parameters $b_{1}$ and $b_{2}$ are the HI linear and quadratic bias parameters, $b_{s^{2}}$ is the tidal bias coefficient, while $b_{\phi}$ and $b_{\phi\delta}$ are the local PNG bias parameters.
$\Omega_{\rm b}$ $\Omega_{\rm c}$ $h$ $A_{\rm s}$ $n_{\rm s}$ $f_{\mathrm{NL}}^{\mathrm{loc}}$
$0.0493$ $0.2644$ $0.6736$ $2.1\times 10^{-9}$ $0.96488$ 0